OpenStudy (anonymous):

if A+B+C= pie,then prove cos2A-cos2B-cos2C=-1+4 cosA sinB sinC

OpenStudy (anonymous):

plsss hlp

OpenStudy (anonymous):

i want to help but i really dont know the ans.

OpenStudy (anonymous):

For answer visit:-

OpenStudy (anonymous):

bt it is another question

OpenStudy (anonymous):

OK

OpenStudy (anonymous):

yeh yeh

OpenStudy (anonymous):

\[A+B+C= \pi \rightarrow A+B = \pi -C \[\sin (A+B)= \sin(\pi-C)= \sin C \] LHS = cos2A-cos2B-cos2C \[= -2 \sin \frac{2A+2B}{2} \sin \frac{2A-2B}{2}-\cos2C\] \[= -2 \sin (A+B) \sin (A-B)-\cos2C\] \[= -2 \sin (\pi -C) \sin (A-B)-\cos2C\] \[= -2 \sin(C) \sin (A-B)-(1-2\sin^2C) = -2 \sin(C) \sin (A-B)-1+2\sin^2C \] \[= 2\sin^2C-2 \sin(C) \sin (A-B)-1 =2sinC [sinC- \sin (A-B)]-1 \] \[= 2sinC[\sin[\pi-(A+B)]- \sin (A-B)]-1 =2sinC [\sin(A+B)- \sin (A-B)]-1\] \[=2sinC [2\cos \frac{(A+B)+(A-B)}{2} \sin \frac{(A+B)-(A-B)}{2}-1\] \[=2sinC [2\cos \frac{A+B+A-B}{2} \sin \frac{A+B-A+B}{2}-1\] \[=4sinC \cos \frac{2A}{2} \sin \frac{2B}{2}-1\] \[=4sinC \cos A \sin B-1\] \[=-1+4sinC \cos A \sin B\] \[=-1+4 \cos A \sin B \sin C \] =RHS Thus LHS=RHS @arkababu

OpenStudy (anonymous):

@arkababu Your problem has been proved. Please check.

OpenStudy (anonymous):

thnkss @dpasingh a lot

OpenStudy (anonymous):

i want to help but i really dont know the ans.